OpenStudy (frostypaz1):

please show your work for this: x^2-8x+16=81

OpenStudy (confluxepic):

Welcome to OpenStudy! @Frostypaz1 Wrong section. :)

Directrix (directrix):

x^2-8x+16 is a perfect square trinomial and follows the factoring pattern: (x - 4)*( x-4) or (x - 4) ^2 81 is a perfect square The equation has become: (x - 4) ^2 = 81 Take square roots of both sides. x - 4 = 9 OR x - 4 = -9 x = 13 OR x = - 5 @Frostypaz1